500 mL of a 0.25 M HCl solution needs to be made from concentrated HCl. What volume of the concentrated solution is needed if its molarity is 12.0 M?



1
Expert's answer
2021-01-21T03:48:16-0500

Proportion 1:

0.25 mol – 1000 mL

x mol – 500 mL

"x = \frac{0.25 \times 500}{1000} = 0.125 \;mol" (quantity of HCl, that must be used)

Proportion 2:

12.0 mol – 1000 mL

0.125 mol – y

"y = \frac{0.125 \times 1000}{12} = 10.41 \;mL" (volume of 12.0 M HCl, that is needed)

Answer: 10.41 mL